# A Look Into News Avoidance Through AWRS: An Avoidance-Aware Recommender System

### v1

@@ Abstract.
Recommender Systems, News Modeling, News Avoidance
1. Introduction
@@
Recommender systems are extensively used to present users with options that align closely with their interests. In the news domain, these systems face unique challenges not typically encountered in other areas, such as the importance of timeliness, novelty, and relevance. These factors create a rapidly changing environment where news consumption patterns shift quickly.
Traditional methods relying solely on user click history and basic profiling have given way to sophisticated techniques that delve deeper into understanding user preferences and behavioral patterns. For instance, the NRMS model (Wu et al., 2019d) uses multi-head self-attention to learn news representations from news titles by modeling the interactions between words. In the user encoder, they learn representations of users from their browsed news and use multi-head self-attention to capture the relatedness between the news. Moreover, the NAML (Wu et al., 2019b) model proposes a neural news recommendation approach which can learn informative representations of users and news by exploiting different kinds of news information such as titles, bodies, and topic categories. Additionally, the LSTUR model (An et al., 2019) learns representations of news from their titles and topic categories and uses an attention network to select important words. In the user encoder, they propose to learn long-term user representations from the embeddings of their IDs.
More recently, other models have tried to increase their performance by focusing on different paradigms. For instance, the GLORY model (Yang et al., 2023) integrates global information with local user interactions to optimize content personalization. Moreover, LANCER (Bae et al., 2023) incorporates the concept of news lifetime to strategically enhance the negative sample space by perceiving that news articles have a finite influence period. Addressing challenges such as the cold-start problem and popularity bias, the PP-Rec model (Qi et al., 2021) utilizes news popularity metrics to enhance recommendation accuracy. In their method, the ranking score for recommending a candidate news to a target user is the combination of a personalized matching score and a news popularity score. Meanwhile, the CAUM model (Qi et al., 2022) leverages candidate-aware self-attention networks to capture global user interests based on candidate news items. They propose a candidate-aware CNN network to incorporate candidate news into local behavior context modeling and learn candidate-aware short-term user interest. Moreover, MANNeR (Iana et al., 2024), a modular framework for flexible multi-aspect (neural) news recommendation that supports ad-hoc customization over individual aspects at inference time, focuses on balancing recommendation performance with diversity across various metrics.
However, one factor that has become increasingly common in today’s digital landscape is the phenomenon of news avoidance. This behavior reflects a deliberate rejection or unintentional neglect of traditional news consumption, often influenced by a preference for alternative media sources (Schrøder, 2019; Villi et al., 2022; Fitzpatrick, 2022). This trend, which can be temporary or selective, underscores a nuanced understanding of how consumers interact with news, shaped by their interests and a growing skepticism towards certain topics. As the demand rises for news that aligns with personal values, especially in a competitive and diverse media landscape, the concept of relevance becomes paramount. Challenges exacerbated by the pandemic, such as news fatigue and distrust in mainstream media, emphasize the importance of recommender systems capable of recognizing and integrating avoidance behaviors.
Despite the aforementioned advancements, the integration of news avoidance behaviors into recommender systems has been overlooked, at least to the best of our knowledge. By engaging domain experts such as journalists and addressing their concerns (Schrøder, 2019; Villi et al., 2022; Fitzpatrick, 2022; Heitz et al., 2022), we explored how to perceive and integrate avoidance strategies within these news recommendation systems. This approach not only enhances the relevance and effectiveness of news recommendations but also acknowledges the evolving dynamics of user engagement with news content in today’s digital landscape.
@@
2. Related Work
2.1. Personalized News Recommender Systems
Neural content-based models have become the leading method for personalized news recommendation, surpassing traditional systems that relied on manual feature engineering (Wu et al., 2023). These Neural News Recommender (NNR) models typically include a news encoder that transforms input features into embeddings and a user encoder that creates user-level representations from clicked news (Wu et al., 2023, 2019d, 2019c; Okura et al., 2017; An et al., 2019; Wu et al., 2022b). The recommendation score is computed by comparing the candidate news embedding against the user embedding (Wang et al., 2018; Wu et al., 2019b). These models are trained using point-wise classification objectives with negative sampling (Huang et al., 2013; Wu et al., 2019d).
Personalized NNR systems focus on matching news content to user preferences, often limiting exposure to diverse viewpoints and creating ”filter bubbles” (Iana et al., 2024; Heitz et al., 2022; Pariser, 2011; Li and Wang, 2019). This leads to homogeneous news consumption, reinforcing users’ initial stances. To address these issues, researchers have explored various methods to diversify recommendations. One approach involves re-ranking personalized recommendation sets (Gharahighehi and Vens, 2021) to increase diversity. Another approach involves multi-task training of NNRs (Moreira et al., 2019; Qi et al., 2022; Wu et al., 2022a; Choi et al., 2022), where the primary personalization objective is coupled with auxiliary objectives that promote aspect-based diversification.
@@ 2.1. Personalized News Recommender Systems
Recent advancements in news recommenders aim to enhance content-based personalization by considering additional aspects: (1) Sentiment - SentiRec (Wu et al., 2020b) recommends news with diverse sentiments. (2) Popularity - PP-Rec (Qi et al., 2021) incorporates news popularity to address cold-start and diversity issues. (3) Global Representations - GLORY (Yang et al., 2023) uses a global news graph and gated graph neural networks to enrich news representations. (4) Recency - LANCER (Bae et al., 2023) introduces the concept of lifetime to improve news recommendation models.
2.2. Open Possibilities and Current Limitations
As previously mentioned, news avoidance has become a growing concern for news platforms, particularly those utilizing recommender systems due to the creation of ”filter bubbles” (Pariser, 2011). We decided to investigate how avoidance affects the behavior of news article clicks. Subsequent sections will illustrate the intrinsic relationship between avoidance, exposure, and the number of clicks a news article receives. This aspect, overlooked in prior studies to the best of our knowledge, can enhance news recommendations by providing insights into the interaction between exposure and avoidance, especially regarding popular items.
In the news domain, the item with the most clicks is often highly exposed and less avoided. We argue that news article avoidance conveys significant information about user preferences. For instance, during an election, election-related articles may be popular but often avoided by those fatigued by political news. If a user reads a political article that is largely avoided by others, it suggests a strong interest in that topic. Conversely, clicking on a popular football article, one that is little avoided, might indicate less about the user’s preference and more about its current popularity. In other words, largely avoided news articles provide more information about user tastes than less avoided ones. This understanding allows us to incorporate such contextual information into recommender systems, improving the understanding of user preferences. The following sections will formally define each of the concepts discussed above.
3. Principal Elements
@@
We identified three principal elements for characterizing a news article: exposure, avoidance, and relevance. To explain these concepts please refer to the Figure 1. In the figure, the black dashed line represents the behaviors dataset, listing news articles (impressions) at each time point $t$. For simplicity, assume each positive news candidate (clicked article) has a fixed number of negative candidates. The off-white square labeled ”exposed news articles” shows all news articles viewed by users at time $t_{1}$. Imagine a news platform where at any time $t$, a certain number of news articles $\mathcal{N}_{t}=\{n_{1},n_{2},...,n_{k}\}$ and users $\mathcal{U}_{t}=\{u_{1},u_{2},...,u_{j}\}$ interact. When user $u_{1}$ views article $n_{1}$, they are ”exposed” to it and can choose to click or not. The box contains all $k$ articles viewed by all $j$ users at time $t$.
Figure 1. A schematic for the explanation of avoidance and exposure.
@@
The light pink rectangle, called the ”impression list,” represents all news articles viewed by user $u$, both positive and negative. An impression log records the articles displayed to a user at a specific time and their click behaviors (Wu et al., 2020a). The set of impressions at time $t$ is termed exposure. Exposure is contingent upon the publication date (time) of news articles before they can be displayed to users. Therefore, at a specific time $t$, certain news articles may be exposed while others are not, simply because they have not yet been published. This concept is illustrated by the arrow at the bottom of the figure, depicting how the impression lists change over time. Finally, we use dark blue and dark pink to indicate whether articles have been clicked or not, respectively.
3.1. Exposure
Definition 3.1 (Number of Exposures - $n_{E}(n,t)$).
@@ 3.1. Exposure
Definition 3.2 (Number of Impressions - $n_{I}(t)$).
@@
Definition 3.3 (Exposure Per Impression - $EPI(n,t)$).
@@
Example: At time $t_{1}$, we know that we have 100 impression lists, and the news article $n_{174}$ has been displayed in 50 of them. Thus, we have $n_{E}(n_{174},t_{1})=50$ and $n_{I}(t_{1})=100$. Hence, the exposure per impression of this news article $n_{174}$ is $EPI(n_{174},t_{1})=\frac{50}{100}=0.5$.
Figure 2. $Av(n,t)\text{ vs. }EPI(n,t)$ for MIND-small.
3.2. Avoidance
Avoidance reflects the deliberate non-engagement with specific news content by readers. This phenomenon may arise from various factors, such as a lack of interest in the topic, discomfort with the content, or distrust towards the source. Understanding the patterns of avoidance is crucial for analyzing user behavior and optimizing content delivery to enhance reader engagement.
Definition 3.4 (Avoidance - $Av(n,t)$).
@@ 3.2. Avoidance
Example: Given that the news article $n_{174}$ at time $t_{1}$ received 20 clicks out of 50 exposures, its avoidance is calculated as $Av(n_{174},t_{1})=1-\frac{n_{clk}(n_{174},t_{1})}{n_{E}(n_{174},t_{1})}=1-\frac{20}{50}=0.6$.
3.2.1. Visualization of Avoidance
We analyzed the correlation among article avoidance, exposure, and the number of clicks for news articles. We plotted $Av(n,t)$ versus $EPI(n,t)$ for the MIND-small dataset on November 9, 2019, at 16:00 and 22:00, as shown in Figure 2. Additionally, we incorporated a color scale to indicate the normalized number of clicks and represented the radius of each news article based on its number of clicks. The graph demonstrates that articles maintaining a balanced level of exposure and low avoidance tend to attract more clicks over time. Articles that show an increase in $EPI$ and a decrease in $Av$ (as indicated by the red arrow) receive the highest number of clicks. For example, article N41881 maintains a stable balance between $EPI$ and $Av$ from 16:00 to 22:00, making it the most clicked news article. In contrast, article N1034 shows a shift in its $EPI$ and $Av$ balance during the same time window. It demonstrates an upward trend in exposure ($\uparrow EPI$) and an increase in avoidance ($\rightarrow Av$) over time, impacting its click rate, which indicates a gradual increase in clicks (larger circle radius). This analysis underscores how avoidance and exposure dynamics influence click rates, offering insights to enhance recommender systems.
@@ 3.2.1. Visualization of Avoidance
(c) Nikkei - Japanese
Figure 3. The graph for $Av(n,t)$ vs $EPI(n,t)$ is computed for the MIND-small (a), Adressa one-week (b), and Nikkei (c) datasets using $D=5$, resulting in 25 distinct regions.
3.3. Relevance
Relevance indicates how well a news article aligns with readers’ interests, but measuring it is complex and requires specific data. Key factors in measuring relevance include: (1) time since publication, (2) news text embedding, (3) exposure per impression $EPI(n,t)$, (4) avoidance $Av(n,t)$, and (5) number of clicks $clk(n,t)$. Each factor is weighted within the model to evaluate relevance, offering a robust framework for assessing the alignment of a news article with user interests. More details on the calculation of relevance can be found in Section 5.3.
4. User Engagement Embedding
To measure the impact of avoidance and exposure on user interaction with a news article, we calculate a metric called user engagement ($\bm{ue}$)—how users interact with a specific news article based on its avoidance and exposure per impression. We divide our graph of $Av(n,t)\text{ vs. }EPI(n,t)$ into $D\times D$ distinct regions. Depending on the values of $Av(n,t)$ and $EPI(n,t)$, a news article falls into different regions. Formally, the user engagement indices are computed as:
(3) $i_{\text{ue}}=D\cdot epi_{\text{idx}}+av_{\text{idx}}$
@@ 4. User Engagement Embedding
(4) $\bm{ue}=\mathbf{W}_{\text{ue}}(i_{\text{ue}})$
where $\mathbf{W}_{\text{ue}}$ is an embedding layer that maps each index to an embedding vector, and $i_{\text{ue}}$ are the computed indices from the above equation 3. The user engagement embedding ($\bm{ue}$) captures the spatial representation based on the quantized values of avoidance and exposure per impression. This embedding is used as input for our models to capture the information conveyed when a specific news article is largely avoided by many users but still clicked by a particular user. The concept behind the user engagement embedding ($\mathbf{ue}$) is to integrate avoidance-awareness into recommender models. Figure 3 illustrates this division for $D=5$, resulting in $D\times D=25$ distinct regions.
5. Methodology
As highlighted in section 2.1, different models focus on various aspects of news recommendations. We introduce a new aspect: Avoidance. Our method, AWRS, improves recommendations by incorporating avoidance-awareness. Unlike traditional methods based only on clicks and popularity, AWRS integrates avoidance and relevance. The conceptual framework is shown in Figure 4.
5.1. Problem Formulation
@@
Figure 4. Overview of our proposed approach for AWRS.
5.2. Document Encoder
@@
Our news encoder, inspired by (Qi et al., 2021, 2022; Wu et al., 2019b), extracts detailed information from text and entities within a news article using a multi-head attention architecture. It incorporates embeddings from the title, category, and entities, generating embeddings with Pretrained Language Models (PLMs) following (Iana et al., 2024). We used RoBERTa Base (Liu et al., 2019) for the MIND-small dataset, NB-BERT Base (Kummervold et al., 2021) for the Adressa one-week dataset, and a finetuned version of the DeBERTa model (He et al., 2021) called Japanese DeBERTa V2111https://huggingface.co/ku-nlp/deberta-v2-base-japanese. For all models, only the last four PLM layers were fine-tuned. The final news embedding n combines $\textbf{n}_{t}$ (title), $\textbf{n}_{c}$ (category), and $\textbf{n}_{e}$ (entity). For the Nikkei and Adressa one-week datasets, only title and category information were used.
5.3. Avoidance-aware Relevance Predictor
Figure 5. (a) Avoidance-aware Relevance Predictor and (b) Avoidance-aware User Encoder Schematics.
@@ 5.3. Avoidance-aware Relevance Predictor
As discussed in Section 3.3, measuring relevance at a given time $t$ involves considering several factors: the content of the article, the time elapsed since publication, avoidance and exposure per impression values, and the number of clicks. Therefore, the Avoidance-aware Relevance Predictor, illustrated in Figure 5 (a), evaluates the relevance of each previously clicked news article $n$. This assessment incorporates the time elapsed, encoded using the Time2Vec (Kazemi et al., 2019) model, resulting in the time elapsed embedding $\bm{t}_{el}$. The news embedding $\bm{n}$, obtained from the news encoder described in Section 5.2, along with the number of clicks at time $t$ denoted as $n_{clk}(n,t)$, and the impact of avoidance and exposure per impression are integrated into the user engagement embedding $\bm{ue}$, as detailed in Section 4. Consequently, the avoidance-aware relevance scores ($\textbf{r}_{aw}$) for each news article are computed as follows:
(5) $\textbf{r}_{aw}=\sigma\left(n_{\textit{clk}}(n,t)\cdot w_{\text{ctr}}+\hat{r}\cdot w_{\hat{r}}\right)$
here, $\hat{r}$ is the weighted sum of the relevance scores influenced by the news article conveyed information, defined as
(6) $\hat{r}=W\cdot\hat{r}_{\text{ic}}+(1-W)\cdot\hat{r}_{\text{tue}}$
here the weight $W$ is computed as $W=\sigma(\Psi_{1}([\textbf{n},\bm{ue},\bm{t}_{el}]))$, where [,,][$\cdot$,$\cdot$,$\cdot$][ ⋅ , ⋅ , ⋅ ] denotes vector concatenation, $\sigma$ is the sigmoid function, and $\Psi_{1}$ is a dense layer. The term $\hat{r}_{\text{ic}}$ represents the relevance score influenced by the news article embedding (n):
(7) $\hat{r}_{\text{ic}}=\Psi_{2}(\textbf{n}),\quad\hat{r}_{\text{tue}}=\Psi_{3}([\bm{ue},\textbf{t}_{el}])$
here the term $\hat{r}_{\text{tue}}$ represents the relevance score influenced by the combination of the user engagement embedding ($\bm{ue}$) and the time elapsed embedding ($\bm{t}_{el}$). Finally, the overall relevance score ($\bm{r}_{aw}$) is obtained by considering both the number of clicks up to time $t$ ($n_{clk}(n,t)$) and the computed relevance score ($\hat{r}$), adjusted by their respective weights ($w_{n_{clk}}$ and $w_{\hat{r}}$) and passed through a sigmoid activation function ($\sigma$) for normalization, as indicated by Equation 5.
@@
5.4. Avoidance-aware User Encoder
The core idea of AWRS is to integrate avoidance-awareness. Leveraging the CAUM model’s architecture (Qi et al., 2022), which facilitates the extraction of how closely candidate news articles relate to previously clicked ones, AWRS enhances the candidate-aware attention network with user engagement embeddings. This enhancement refines the comprehension of user preferences by examining the categories of news articles users interact with, as well as their levels of exposure and avoidance. This approach yields a nuanced depiction of user interests.
5.4.1. Historical and Candidate News Vector Concatenation
@@
For each historical and candidate news item, we enhance its representation by appending the corresponding user engagement embeddings. This generates enriched feature vectors for a more tailored recommendation approach. Given a set $\mathbf{N}_{c}$ of candidate news articles and a set $\mathbf{H}_{u}$ of historical clicks for a user $u$, we concatenate the user engagement embedding $\bm{ue}$ for each item. Let $\mathscr{H}$ and $\mathscr{N}$ be the sets with concatenated user engagement embeddings. Formally, for each $i$-th item $\bm{h}_{i}=[\textbf{h}_{u_{i}},\bm{ue}_{h_{i}}]$ and $\bm{n}_{i}=[\mathbf{n}_{c_{i}},\bm{ue}_{c_{i}}]$. Where $\bm{ue}_{h_{i}}$ and $\bm{ue}_{c_{i}}$ are the user engagement embeddings for the $i$-th historical and candidate news items, respectively, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes vector concatenation. Thus, we have the sets:
(8) $\mathscr{H}=[\bm{h}_{1},...,\bm{h}_{M}],\quad\mathscr{N}=[\bm{n}_{c1},...,\bm{n}_{cL}]$
5.4.2. Avoidance-aware Self-Attention Layer
@@
Given the history of clicked news articles $\mathscr{H}_{u}^{M}$, where $M$ represents the total number of articles previously clicked by the user $u$, we use multiple self-attention heads to assess similarities and extract relatedness information between the $i$-th and $j$-th historical clicks, $\textbf{h}_{i}$ and $\textbf{h}_{j}$, respectively:
(9) $\hat{r}^{k}_{i,j}=\textbf{q}^{T}_{i}\textbf{W}^{r}_{k}\bm{h}_{j},\quad\textbf{q}^{T}_{i}=\textbf{Q}_{u}\bm{h}_{i}$
here, $\hat{r}^{k}_{i,j}$ is the attention score from the $k$-th head, with $\textbf{Q}_{u}$ as the projection matrix and $\textbf{W}^{r}_{k}$ as the parameters of the $k$-th attention head. We adaptively select significant long-range relatedness metrics to model user interest in the candidate news $\mathscr{N}$ based on their contextual relevance
(10) $r^{k}_{i,j}=\hat{r}^{k}_{i,j}+\textbf{q}^{T}_{c}\textbf{W}^{r}_{k}\bm{h}_{j}$

### v2

@@ Abstract
Keywords: Recommender Systems, News Modeling, News Avoidance
1 Introduction.
@@
Recommender systems are extensively used to present users with options that align closely with their interests. In the news domain, these systems face unique challenges not typically encountered in other areas, such as the importance of timeliness, novelty, and relevance.
Traditional methods based on user click history have evolved into advanced techniques that better capture user preferences and behaviors. The NRMS model [14], for example, uses multi-head self-attention to learn news representations from titles and to capture user preferences from their browsing history. The NAML model [13] utilizes convolutional neural networks (CNNs) and enriches user and news representations by leveraging various types of news information, such as titles, bodies, and categories. The LSTUR model [15] focuses on both short- and long-term user representations and selects important words through an attention network. Recent models have further advanced performance through new approaches. The GLORY model [21] integrates global information with local user interactions for better personalization. LANCER [43] uses the concept of news lifetime to enhance negative sampling by considering the finite influence period of news. The PP-Rec model [27] tackles cold-start and popularity bias by combining personalized and popularity-based ranking scores. CAUM [22] leverages candidate-aware self-attention networks to capture global user interests and incorporates candidate news into local context modeling.
A growing trend in today’s digital landscape is news avoidance, where users deliberately reject or neglect certain news, influenced by a preference for alternative media [2, 1, 3]. This behavior, often temporary or selective, highlights the need for recommender systems to integrate avoidance patterns. According to [3], mental health concerns are identified as the primary factor driving the increase in news avoidance, with growing distrust in mainstream media also being a significant contributor. Although skepticism toward journalism is not a new issue, it was amplified during the pandemic by the influence of anti-vaccination advocates and conspiracy theorists. Despite advancements in news modeling for recommendation systems, the integration of news avoidance behaviors remains largely unexplored.
@@
2 Related Work.
Neural networks models have emerged as the leading approach for personalized news recommendations, leveraging news and user encoders to compute recommendation scores by comparing candidate news embeddings with user embeddings [39, 14, 16, 24, 44, 47, 48, 22].
@@
Recent advancements in news recommender systems have primarily focused on four key aspects: (1) Sentiment – SentiRec [23], which recommends news with diverse sentiment profiles; (2) Popularity – PP-Rec [27], which leverages news popularity to address cold-start issues and enhance diversity; (3) Global Representations – GLORY [21], which enriches news representations by utilizing a global news graph and gated graph neural networks; and (4) Recency – LANCER [43], which incorporates the concept of news lifetime to improve recommendation accuracy.
In this work, we introduce a novel perspective on news recommendation by integrating avoidance awareness, aiming to enhance the understanding of user engagement patterns and preferences.
3 Principal Elements.
Figure 1: Avoidance and exposure explanation diagram
@@
We identified three principal elements for characterizing a news article: exposure, avoidance, and relevance. Before defining these concepts, please refer to Figure 1, which is needed to illustrate some terms. In the figure, the black dashed line represents the behaviors dataset, listing news articles (impressions) at each time point $t$. For simplicity, assume that each clicked news article (positive candidate) has a fixed number of non-clicked articles (negative candidates). The section labeled “exposed news articles” includes all the news articles viewed by users at time $t_{1}$. Consider a news platform where, at any given time $t$, a set of news articles $\mathcal{N}_{t}=\{n_{1},n_{2},\ldots,n_{k}\}$ and a group of users $\mathcal{U}_{t}=\{u_{1},u_{2},\ldots,u_{n}\}$ interact. When a user $u_{1}$ views an article $n_{1}$, they are “exposed” to it and can choose to click on it or not. This section contains all the articles $\mathcal{N}$ viewed by all users $\mathcal{U}$ at a specific time $t$.
@@
The “impression list” represents all news articles shown to a user $u$, including both those that were clicked and those that were not. An impression log records the articles displayed to a user at a specific time and their click behaviors [31]. The set of impressions at time $t$ is referred to as exposure. Exposure depends on the publication date (time) of news articles; only articles published before the exposure time can be displayed to users. Therefore, at a specific time $t$, some news articles may be exposed while others are not, simply because they have not yet been published. This concept is illustrated by the arrow at the bottom of the figure, which shows how the impression lists change over time.
3.1 Exposure.
Definition 1 (Number of Exposures - $n_{E}(n,t)$).
@@
Definition 2 (Number of Impressions - $n_{I}(t)$).
@@
Definition 3 (Exposure Per Impression - $EPI(n,t)$).
@@
Example: At time $t_{1}$, the news article $n_{174}$ appears in 50 out of 100 impression lists, giving $n_{E}(n_{174},t_{1})=50$ and $n_{I}(t_{1})=100$. Thus, its exposure per impression is $EPI(n_{174},t_{1})=\frac{50}{100}=0.5$.
3.2 Avoidance.
Avoidance represents a reader’s decision to disengage from specific news content due to factors such as disinterest, discomfort, or distrust. Analyzing avoidance patterns is key to understanding user behavior and optimizing content recommendations.
Definition 4 (Avoidance - $Av(n,t)$).
@@
Example: The avoidance of news article $n_{174}$ at time $t_{1}$, with 20 clicks from 50 exposures, is $Av(n_{174},t_{1})=1-\frac{n_{clk}(n_{174},t_{1})}{n_{E}(n_{174},t_{1})}=1-\frac{20}{50}=0.6$.
3.2.1 Consequences of Avoiding News Articles.
Consider a large set of users exposed to a news article $n_{174}$ who chose not to click on it, either deliberately or unintentionally, indicating a lack of interest. Now, suppose a specific user, $u_{5}$, does click on it. Since most users avoided the article, $u_{5}$’s engagement reveals deeper insights into their preferences compared to interactions with widely popular articles. Thus, engaging with avoided articles can better reflect a user’s unique interests, while avoidance serves as a measure of disinterest or resistance to a topic.
3.2.2 Visualization of Avoidance.
We analyzed the correlation between news article avoidance, exposure, and the number of clicks per article. Figure 2 shows $Av(n,t)$ versus $EPI(n,t)$ for the MIND-small [31] dataset on November 9, 2019, from 00:00 AM to 08:00 AM. The color scale represents the normalized click count, and the radius of each article reflects its number of clicks, scaled by a factor of three for better visualization.
Note how the overall number of clicks increases over time, as expected, since more users log into the platform. Additionally, we have highlighted two specific articles: one with a black square and the other with a black circle. Notably, the article marked with the black circle consistently received more clicks over time compared to all other articles. This article maintains a high $EPI$ and low $Av$, with $Av(n,t)$ remaining below 0.8 (red dashed line) at all times, maintaining the highest normalized number of clicks.
In contrast, the article marked with the black square starts with a green color, but as time progresses, its avoidance increases while its exposure per impression decreases. As a result, its relative number of clicks gradually declines, with the color shifting from green to blue. This suggests that articles exhibiting an increase in $EPI$ and a decrease in $Av$ (as indicated by the red arrow in the bottom-right chart) tend to receive the highest number of clicks.
Figure 2: $Av(n,t)\text{ vs. }EPI(n,t)$ for MIND-small [31] at 2019-11-09 from 00:00 AM to 08:00 AM.
This analysis demonstrates how the dynamics of avoidance and exposure influence click behavior. For a detailed hour-by-hour visualization of avoidance changes, please refer to Appendix B.
@@
(c) Nikkei one-week - Japanese
Figure 3: The graph of $Av(n,t)$ versus $EPI(n,t)$ is generated for the MIND-small [31] (a), Adressa one-week [30] (b), and Nikkei one-week (c) datasets, using $D=5$, which results in 25 distinct regions.
Figure 4: Overview of AWRS.
3.3 Relevance.
Indicates how well a news article matches reader interests, requiring specific data like (1) time since publication, (2) text embedding, (3) exposure per impression $EPI(n,t)$, (4) avoidance $Av(n,t)$, and (5) clicks $clk(n,t)$. The model weighs these factors to assess relevance, aligning news with user interests. See Section 5.3 for details on relevance calculation.
4 User Engagement Embedding.
To measure the impact of avoidance and exposure on user interaction with a news article, we calculate a metric called user engagement ($\bm{ue}$)—how users interact with a specific news article based on its avoidance and exposure per impression. We divide our graph of $Av(n,t)\text{ vs. }EPI(n,t)$ into $D\times D$ distinct regions. Depending on the values of $Av(n,t)$ and $EPI(n,t)$, a news article falls into different regions.
Formally, the user engagement indices are computed as:
(4.3) $i_{\text{ue}}=D\cdot epi_{\text{idx}}+av_{\text{idx}}$
@@
(4.4) $\bm{ue}=\mathbf{W}_{\text{ue}}(i_{\text{ue}})$
where $\mathbf{W}_{\text{ue}}$ is an embedding layer that maps each index to an embedding vector, and $i_{\text{ue}}$ are the computed indices from the above equation 4.3. The user engagement embedding ($\bm{ue}$) captures the spatial representation based on the quantized values of avoidance and exposure per impression. This embedding is used as input for our models to capture the information conveyed when a specific news article is largely avoided by many users but still clicked by a particular user. The concept behind the user engagement embedding ($\mathbf{ue}$) is to integrate avoidance-awareness into recommender models. Figure 3 illustrates this division for $D=5$, resulting in $D\times D=25$ distinct regions.
5 Methodology.
Section 2 discusses various aspects of news recommendations. We introduce a new aspect: Avoidance. Our method, AWRS, unlike traditional methods based only on clicks and popularity, integrates avoidance awareness. The conceptual framework is shown in Figure 4.
5.1 Problem Formulation.
@@
5.2 Document Encoder.
@@
Our news encoder, inspired by [27, 22, 13], employs a multi-head attention mechanism to extract information from embeddings generated by Pretrained Language Models (PLMs) [24]. We use RoBERTa Base [35] for MIND-small, NB-BERT Base [38] for Adressa one-week, and fine-tuned Japanese DeBERTa V2111https://huggingface.co/ku-nlp/deberta-v2-base-japanese for Nikkei one-week, with only the last four PLM layers fine-tuned. The final news embedding n combines title ($\textbf{n}_{t}$), category ($\textbf{n}_{c}$), and entity ($\textbf{n}_{e}$) embeddings, while for Nikkei one-week and Adressa one-week, only title and category embeddings are used.
5.3 Avoidance-aware Relevance Predictor.
@@
As discussed in Section 3.3, measuring relevance at a given time $t$ involves considering several factors: the content of the article, the time elapsed since publication, avoidance and exposure per impression values, and the number of clicks. Therefore, the Avoidance-aware Relevance Predictor, illustrated in Figure 5 (a), evaluates the relevance of each previously clicked news article $n$.
This assessment incorporates the time elapsed, encoded using the Time2Vec [51] model, resulting in the time elapsed embedding $\bm{t}_{el}$. The news embedding $\bm{n}$, obtained from the news encoder described in Section 5.2, along with the number of clicks at time $t$ denoted as $n_{clk}(n,t)$, and the impact of avoidance and exposure per impression are integrated into the user engagement embedding $\bm{ue}$, as detailed in Section 4. Consequently, the avoidance-aware relevance scores ($\textbf{r}_{aw}$) for each news article are computed as follows:
@@
Figure 5: (a) Avoidance-aware Relevance Predictor and (b) Avoidance-aware User Encoder Schematics.
(5.5) $\textbf{r}_{aw}=\sigma\left(n_{\textit{clk}}(n,t)\cdot w_{\text{ctr}}+\hat{r}\cdot w_{\hat{r}}\right)$
here, $\hat{r}$ is the weighted sum of the relevance scores influenced by the news article conveyed information, defined as $\hat{r}=W\cdot\hat{r}_{\text{ic}}+(1-W)\cdot\hat{r}_{\text{tue}}$. Where the weight $W$ is computed as $W=\sigma(\Psi_{1}([\textbf{n},\bm{ue},\bm{t}_{el}]))$, where [,,][$\cdot$,$\cdot$,$\cdot$][ ⋅ , ⋅ , ⋅ ] denotes vector concatenation, $\sigma$ is the sigmoid function, and $\Psi_{1}$ is a dense layer.
The term $\hat{r}_{\text{ic}}$ represents the relevance score influenced by the news article embedding (n):
(5.6) $\hat{r}_{\text{ic}}=\Psi_{2}(\textbf{n}),\quad\hat{r}_{\text{tue}}=\Psi_{3}([\bm{ue},\textbf{t}_{el}])$
here the term $\hat{r}_{\text{tue}}$ represents the relevance score influenced by the combination of the user engagement embedding ($\bm{ue}$) and the time elapsed embedding ($\bm{t}_{el}$).
Finally, the overall relevance score ($\bm{r}_{aw}$) is obtained by considering both the number of clicks up to time $t$ ($n_{clk}(n,t)$) and the computed relevance score ($\hat{r}$), adjusted by their respective weights ($w_{n_{clk}}$ and $w_{\hat{r}}$) and passed through a sigmoid activation function ($\sigma$) for normalization, as indicated by Equation 5.5.
5.4 Avoidance-aware User Encoder.
The core concept of AWRS is avoidance-awareness. Building on CAUM’s architecture [22], which focuses on capturing the relationship between candidate and previously clicked news, AWRS enhances candidate-aware attention with user engagement embeddings. This addition provides a deeper understanding of user preferences by analyzing news categories, exposure, and avoidance, resulting in a more refined user interest representation.
5.4.1 Historical and Candidate News Vector Concatenation.
@@
For each historical and candidate news item, we enrich its representation by appending the corresponding user engagement embedding, resulting in enhanced feature vectors that support a more personalized recommendation approach. Given a set $\mathbf{N}_{c}$ of candidate news articles and a set $\mathbf{H}_{u}$ of historical clicks for a user $u$, we concatenate the respective user engagement embedding $\bm{ue}$ for each item.
Let $\mathscr{H}$ and $\mathscr{N}$ denote the sets of historical and candidate news articles with their user engagement embeddings concatenated. Formally, for each $i$-th item:
(5.7) $\bm{h}_{i}=[\textbf{h}_{u_{i}},\bm{ue}_{h_{i}}]\quad\text{and}\quad\bm{n}_{i}=[\mathbf{n}_{c_{i}},\bm{ue}_{c_{i}}]$
where $\bm{ue}_{h_{i}}$ and $\bm{ue}_{c_{i}}$ represent the user engagement embeddings for the $i$-th historical and candidate news items, respectively, and the operator [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes vector concatenation. Thus, the resulting sets are defined as:
(5.8) $\mathscr{H}=[\bm{h}_{1},\dots,\bm{h}_{M}]\quad\text{and}\quad\mathscr{N}=[\bm{n}_{1},\dots,\bm{n}_{L}]$
which now incorporate avoidance information into their representations for each historical and candidate item.
5.4.2 Avoidance-aware Self-Attention Layer.
@@
Given the user’s click history $\mathscr{H}_{u}^{M}$, where $M$ is the total number of previously clicked articles, we employ multiple self-attention heads to capture the relatedness between the $i$-th and $j$-th historical clicks, $\textbf{h}_{i}$ and $\textbf{h}_{j}$, respectively. Specifically, we compute:
(5.9) $\hat{r}^{k}_{i,j}=\textbf{q}^{T}_{i}\textbf{W}^{r}_{k}\bm{h}_{j},\quad\textbf{q}^{T}_{i}=\textbf{Q}_{u}\bm{h}_{i}$
where $\hat{r}^{k}_{i,j}$ represents the attention score from the $k$-th head, $\textbf{Q}_{u}$ is the projection matrix, and $\textbf{W}^{r}_{k}$ denotes the parameters of the $k$-th attention head. This approach captures relatedness not only from the news content but also from the user engagement information.